OpenStudy (anonymous):

Calculus: Find equations of the tangents to the curve x=(3t^2)+1, y= (2t^3)+1 that pass through the point (4,3). (I know that I need to find dy/dx, but when I solve for t, since i have 2 equations, I have 2 ts, and don't know which one to use.

OpenStudy (anonymous):

dy/dx = (dy/dt)*(dt/dx)

OpenStudy (anonymous):

The point (4,3) corresponds to t=1

OpenStudy (anonymous):

What is \[ \frac {dy}{dt}= ?\\ \frac {dx}{dt} = ?\\ \frac {\frac {dy}{dt}} {\frac {dx}{dt}}=? \]

OpenStudy (anonymous):

I will do one for you \[ y= 2t^3 +1\\ \frac {dy}{dt} = 6 t^2 \]

OpenStudy (anonymous):

then \[ x=3t^2+1\\ \frac { dx}{dt}= 6t \]

OpenStudy (anonymous):

The ratio of the two is \[ \frac {6 t^2}{6t}=t \] so \[\frac {dy}{dx}=t=1 \] at the point. The slope of the tangent is 1 at the point (4,3)\\ the equation of the tangent y-3 = 1(x-4) y= x-4 +3 = x -1